The new drill – University of Copenhagen

Forward this page to a friend Resize Print Bookmark and Share

EastGrip - The East Greenland Ice-core Project > Field Diaries Folder > UK diaries 2017 > The new drill

25 May 2017

Thursday 25th of May 2017

Scene from the assembly of warm laboratories in the science trench.

The new drilling electronics is temperamental.

Our drillers, software guy and electrician play a game these days. The new drill electronics is re-programmed and re-wired, and sent down with the drill for testing. Then an issue arises, and the electronics is replaced with the less advanced but reliable HT electronics. The drillers then make some good, ice core producing runs, while the new electronics is under modification. It's slow progress, but our feeling is that eventually we will reach the goal.

Today, another issue was solved: The suction hose for ice chips got clogged several times as the ice chips packed and froze in the hose. A heating cable on the suction hose seems to have helped.

What we did today:

  1. Check of fuel pump setting on main generator. Camp was on back-up power for a short while.
  2. Drilling 5 runs.
  3. Logging ice cores.
  4. Working on infrastructure in science trench.
  5. Outfitting warm laboratories.
  6. Testing radar.
  7. Site for future cooks freezer has been selected and marked.
  8. Laying cables in tunnel between main generator and carpenters garage.
  9. German snow sampling.
  10. Japanese aerosol sampling and snow experiments.
  11. Water vapour sampling, methane sampling, boundary layer and meteorology station working.

Ad. 2: Drilllers report: Test run with new electronics and RECAP drill, failed after 30 cm. Continued liquid drilling, 4 runs with HT electronics and RECAP drill, 1.7 m to 2.0 m cores. Modified plumbing for chip suction. Drillers depth: 123.46m

Ad.3: Logging depth, start: 126.58 m. Logging depth, end: 132.96 m.

Ad.4: Horizontal saw installed. Tables prepared for bandsaws. Tables set up around warm laboratories.

Weather today: Sunshine. Temp. - 14°C to - 24°C. Wind: 5 - 10 kt from W. Visibility: To horizon.

FL, J.P. Steffensen

Thursday 25th of May 2017

Scene 2 from the assembly of warm laboratories in the science trench.

← Previous entry   Next entry →